OpenStudy (anonymous):

prove: [(3x+2)/x-1]-6/5=0

OpenStudy (zarkon):

prove?

OpenStudy (anonymous):

solve.. yes..

OpenStudy (anonymous):

please?

OpenStudy (zarkon):

is it \[\frac{3x+2}{x-1}-\frac{6}{5}=0\] or \[\frac{3x+2}{x}-1-\frac{6}{5}=0\]

OpenStudy (anonymous):

the first one..

OpenStudy (zarkon):

\[\frac{3x+2}{x-1}-\frac{6}{5}=0\] \[\frac{3x+2}{x-1}=\frac{6}{5}\] \[3x+2=\frac{6}{5}(x-1)\] multiply through by 5 \[15x+10=6(x-1)\] \[15x+10=6x-6)\] \[9x=-16\] \[x=-\frac{16}{9}\]

OpenStudy (anonymous):

thank you so much